OpenStudy (anonymous):

Analyze the graph of the following function as follows: (a) Find the x- and y-intercepts. (b) Determine the end behavior: find the power function that the graph of f resembles for large values of |x|. (c) Find the maximum number of turning points. (d) Graph the function Please show all of your work. f(x)=x^2(x+2)

OpenStudy (anonymous):

is this \[f(x)=x^2(x+2)=x^3+2x^2\]?

OpenStudy (anonymous):

zeros are \[x=0,x=-2\]

OpenStudy (anonymous):

end behaviour looks like x^3 so goes to infinity as x does, and minus infinity as x goes to minus infinity

OpenStudy (anonymous):

degree is 3 so it can have at most two turning points

OpenStudy (anonymous):

and y - intercept is what you get when x = 0, namely 0

OpenStudy (anonymous):

Given Equation: \[x ^{2}(x+2)\] Which is, \[x ^{3}+2x ^{2}\] So our equation becomes, \[F(x)=x ^{3}+2x ^{2}\] Or, \[Y=x ^{3}+2x ^{2}\] 1) Finding the x-intercepts and y-intercepts. Now, x-intercept: The point at which the graph cuts the x-axis. To find the x-intercept, Take Y=0 So, \[0=x ^{3}+2x ^{2}\] Take 2x^2 to other side, \[-2x ^{2}=x ^{3}\] Cancelling x^2 and x^3 from both sides, we are left with: \[-2=x\] So the x-intercept is -2. Finding the y-intercept: Y-intercept: The point at which the graph cuts the y-axis. So, to find the y-intercept, take x=0 So, F(x)=F(0): \[(0)^{3}+2(0)^{2}\] Which ultimately gets us 0. So the y-intercept is 0. So we get the coordinates of the point as (-2,0) The graph should be something like this... 2.I didn't understand sorry... 3. There are no turning points, it is a straight line, so 0 turning points. 4.Graph attached. Please do ask if you need any more help! :D

OpenStudy (asnaseer):

@satellite73 has given you all the required information - you just need to graph it now.

OpenStudy (anonymous):

Correction: there can be 3 turning points like Satellite73 said, I didn't notice the degree earlier sorry.

OpenStudy (anonymous):

how is the graph supposed to go can someone show me please?